# https://neu-vi.github.io/HOI-Diff/

### v2

@@ 3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [54] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Fig. 4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
@@
Affordance estimation in 3D point clouds itself is a notably challenging problem [30, 7, 22, 16, 29, 20, 21], especially in the context of 3D HOI generation involving textual prompt. In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios. It can effectively model common interactions such as grasping an object with both hands, sitting actions involving the pelvis and back, or lifting with a single hand. We use binary contact labels to determine which joints are in contact with the object. Subsequently, we predict eight corresponding contact points on the object surface, identified as the points closest to the selected body joints. Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
@@
Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and HOI-DM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$. The model is designed to predict both contact probabilities and contact positions. Subsequently, dynamic selection of contacting body joints is performed by considering predicted probabilities over a specific threshold $\tau$ (set to be 0.6). The corresponding contact points on the object are then determined based on the selected joints. APDM works similar to the diffusion denoising process described in Eq.(1). Besides, we utilize a large language model (ChatGPT) to determine whether the object state ${\bm{y}}_{0}^{s}\in\{0,1\}$ should be set to static (${\bm{y}}_{0}^{s}=1$) based on the textual description, which can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
@@ 3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [8] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Algorithm 1 Affordance-guided Interaction Correction 1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$. 2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$ 3:$K$ = 1 4:for all $t$ from $T$ to $1$ do 5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$ 6:     if $t=1$ then 7:         $K=100$ 8:     end if 9:     for all $k$ from $K$ to $1$ do # Separately perturb 10:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$ 11:     end for 12:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$ 13:end for 14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [53, 19, 23],
@@
Here $\tau_{1}$ and $\tau_{2}$ are different strengths to control the guidance for human and object motion, respectively. Due to the sparseness of object motion features, we assign a larger value to $\tau_{2}$ compared to $\tau_{1}$. This applies greater strength to perturb object motion, facilitating feasible corrections for contacting joints. During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step, as illustrated in Algorithm 1.
@@
$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
@@
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in the supplementary material.
@@
$G=G_{cont}+\alpha G_{sta}+\beta G_{smo},$ (7)
where $\alpha$ and $\beta$ are weights for balance.
@@ 4.1 Setup
Dataset. Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [3]. We then use ChatGPT to assist in generating three text descriptions for each HOI sequence, increasing the diversity of the data. Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects. It provides the human SMPL-H representation [27], the object mesh, as well as its 6DoF pose information in each HOI sequence. To ensure consistency in our approach, we follow the processing method used in HumanML3D [11] to extract representations for 22 body joints. All the models are trained to generate $L=196$ frames in our experiments. In the end, we have 1451 3D HOI sequences along with textual descriptions to train and evaluate our proposed approach. We follow the official train/test split on BEHAVE. We provide more details of the dataset, our annotation process, and annotated textual examples in the supplementary material.
In addition, we evaluate our approach on OMOMO dataset [25]. OMOMO focuses on full-body manipulation with hands. It consists of human-object interaction motion for 15 objects in daily life, with a total duration of approximately 10 hours. It provides text descriptions for each interaction motion. We utilize their object split strategy for both training and evaluation, ensuring the objects between the training and testing sets are different. Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset. More details of the OMOMO dataset are in the supplementary material.
Evaluation metrics. We first assess different models for human motion generation using standard metrics as introduced by [11], namely Fréchet Inception Distance (FID), R-Precision, and Diversity. FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder. R-Precision gauges the relevance between generated motions and their corresponding text prompts. Diversity evaluates the range of variation in the generated motions. Additionally, we compute the Foot Skating Ratio to measure the proportion of frames exhibiting foot skid over a threshold (2.5 cm) during ground contact (foot height $<$ 5 cm).
To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points. Ideally, we should develop similar metrics, e.g., FID, to evaluate the stochastic HOI generation. However, due to the limited data available in BEHAVE [3], training a motion encoder would produce biased evaluation results. To mitigate this issue, we resort to user studies to quantify the effectiveness of different models. Details will be introduced later.

### v3

@@ 3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, DBDM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [xu2023interdiff] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g, “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Figure 4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the DBDM and thus can recover from the potential errors in DBDM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
@@
Affordance estimation in 3D point clouds itself is a notably challenging problem [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive], especially in the context of 3D HOI generation involving textual prompt. In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios. It can effectively model common interactions such as grasping an object with both hands, sitting actions involving the pelvis and back, or lifting with a single hand. We use binary contact labels to determine which joints are in contact with the object. Subsequently, we predict eight corresponding contact points on the object surface, identified as the points closest to the selected body joints. Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
@@
Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and DBDM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$. The model is designed to predict both contact probabilities and contact positions. Subsequently, dynamic selection of contacting body joints is performed by considering predicted probabilities over a specific threshold $\tau$ (set to be 0.6). The corresponding contact points on the object are then determined based on the selected joints. APDM works similar to the diffusion denoising process described in Eq.(1). Besides, we utilize a large language model (ChatGPT) to determine whether the object state ${\bm{y}}_{0}^{s}\in\{0,1\}$ should be set to static (${\bm{y}}_{0}^{s}=1$) based on the textual description, which can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$. More implementation details are in Appendix A.2.
@@ 3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [dhariwal2021diffusion] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [xie2023omnicontrol, karunratanakul2023guided],
@@
Here $\tau_{1}$ and $\tau_{2}$ are different strengths to control the guidance for human and object motion, respectively. Due to the sparseness of object motion features, we assign a larger value to $\tau_{2}$ compared to $\tau_{1}$. This applies greater strength to perturb object motion, facilitating feasible corrections for contacting joints. During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step. The details are illustrated in Algorithm 1 of Appendix.
@@
$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{\mu}}^{o}_{t},{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
@@
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in Appendix A.3.
@@
$G=G_{con}+\alpha G_{sta}+\beta G_{smo},$ (7)
where $\alpha=500$ and $\beta=100$ are weights for balance.
@@ 4.1 Setup
Dataset. Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [bhatnagar22behave]. We then use GPT-3.5 [chatgpt] to rephrase and generate three text descriptions for each HOI sequence, increasing the diversity of the data. Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects. It provides the human SMPL-H representation [SMPL:2015], the object mesh, as well as its 6DoF pose information in each HOI sequence. To ensure consistency in our approach, we follow the processing method used in HumanML3D [Guo_2022_CVPR] to extract representations for 22 body joints. All the models are trained to generate $L=196$ frames in our experiments. In the end, we have 1451 3D HOI sequences along with textual descriptions to train and evaluate our proposed approach. We follow the official train/test split on BEHAVE. We provide more details of the dataset and annotation process in Appendix I.
In addition, we evaluate our approach on OMOMO dataset [li2023object]. OMOMO focuses on full-body manipulation with hands. It consists of human-object interaction motion for 15 objects in daily life, with a total duration of approximately 10 hours. It provides text descriptions for each interaction motion. We utilize their object split strategy for both training and evaluation, ensuring the objects between the training and testing sets are different. Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset. More details are in Appendix J.
Evaluation metrics. We first assess different models for human motion generation using standard metrics as introduced by [Guo_2022_CVPR], namely Fréchet Inception Distance (FID), R-Precision, and Diversity. FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder. R-Precision gauges the relevance between generated motions and their corresponding text prompts. Diversity evaluates the range of variation in the generated motions. Additionally, we compute the Foot Skating Ratio to measure the proportion of frames exhibiting foot skid over a threshold (2.5 cm) during ground contact (foot height $<$ 5 cm).
To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points. Ideally, we should develop similar metrics, e.g, FID, to evaluate the stochastic HOI generation. However, due to the limited data available in BEHAVE [bhatnagar22behave], training a motion encoder would produce biased evaluation results. To mitigate this issue, we resort to user studies to quantify the effectiveness of different models. Details will be introduced later.
@@ 4.2 Comparisons with Existing Methods
Baselines. Our work introduces a novel 3D HOIs generation task not addressed by existing text-to-motion methods, which focus exclusively on human motion generation without accounting for human-object interactions. To compare with existing works, we mainly focus on evaluating human motion generation. We then design different variants of our models for comparing 3D HOIs generation. Specifically, we adopt the prominent text-to-motion methods MDM [tevet2023human] and PriorMDM* [shafir2023human] with the following settings. (a) MDM: In this setup, we finetune the original MDM model [tevet2023human] on the BEHAVE dataset [bhatnagar22behave] without object motion. (b) MDM*: This variant involves adapting the input and output layers’ dimensions of the MDM model [tevet2023human] to accommodate the input of 3D HOI sequences. This adjustment allows for the simultaneous learning of both human and object motions within a singular, integrated model. (c) PriorMDM* [shafir2023human]: We adapt the ComMDM architecture proposed in [shafir2023human], originally designed for two-person motion generation, to suit our needs for HOIs synthesis by modifying one of its two branches for object motion generation. (d) InterDiff [xu2023interdiff]: While InterDiff is not designed for text-driven synthesis of 3D HOI, we added text conditioning to InterDiff as the baseline. More details are in Appendix C.
BEHAVE OMOMO
Method FID
R-precision
Diversity
Contact
Pene
Foot Skate
FID
R-precision
Diversity
Contact
Pene
Foot Skate
Ratio $\downarrow$
Real 0.04 0.86 12.48 - - - 0.57 0.63 9.98 - - -
MDM 6.77 0.34 10.81 - - - 12.28 0.23 5.56 - - -
MDM* 4.25 0.38 11.23 0.448 0.52 0.190 10.37 0.21 6.04 0.768 0.41 0.191
PriorMDM* 4.54 0.30 10.03 0.416 0.57 0.270 9.87 0.25 6.34 0.523 0.38 0.344
InterDiff 8.58 0.26 10.75 0.506 0.42 0.218 14.27 0.17 5.69 0.906 0.32 0.239
Ours 1.62 0.46 12.02 0.347 0.51 0.182 8.76 0.31 8.13 0.326 0.39 0.141
Table 1: Quantitative results on the BEHAVE and OMOMO dataset. We compare our method with baselines adapted from existing models. MDM: fine-tune the original MDM [tevet2023human] on the BEHAVE dataset without object motion. MDM*: adapting the input and output layers’ dimensions of the MDM to accommodate both human and object motions. PriorMDM*: We adapt the ComMDM architecture proposed in shafir2023human. InterDiff: We add a CLIP encoder in xu2023interdiff to support our task. The right arrow $\rightarrow$ means closer to real data is better.
Figure 5: Qualitative comparisons of our approach and baselines on BEHAVE dataset. The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4. This contrasts with the baselines, which fail to achieve a similar level of realism and contact plausibility in the interactions. As an additional visual aid, the mesh color gradually darkens over time to represent progression. (Best viewed in color.)
Quantitative Results. Table 4.2-left reports the quantitative results on BEHAVE dataset [bhatnagar22behave]. Compared with the baseline methods, our full method achieves the best performance. Specifically, it achieves state-of-the-art results in both FID, R-precision, and Diversity, underscoring its ability to generate high-quality human motions in the context of coherently interacting with objects. The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects. Table 4.2-right presents the quantitative results on the OMOMO dataset. We used the train/test split of the OMOMO dataset to evaluate the model’s inference capacity on unseen objects, including the small table, white chair, suitcase, and tripod. Our method consistently outperforms other baselines by a considerable margin across all metrics. Notably, due to the distinctiveness of objects in the training and testing sets, the results indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models. We also provide user study results, please refer to Appendix G for details.
Qualitative Results. We showcase qualitative comparisons, rendered with SMPL [SMPL:2015] shapes, between our approach and the baseline methods in Figure 5. It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g, the toolbox in Figure 5 (b)). Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g, the chair in Figure 5 (e)). In stark contrast, our approach excels in creating visually appealing and realistic HOIs. Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g, raising the trash bin with the right hand in Figure 5 (a)). For the same object, our method can generate diverse HOIs using different body parts and contact points, as shown in Figure 16 in Appendix.
Supplementary Material
Appendix A Additional Details of Methodology
In Sec. 3 of our main paper, we presented the foundational design of each key component in our HOI-Diff pipeline. Here, we delve into an elaborate explanation of model architecture, learning objectives and additional details associated with each crucial component.
A.1 Dual-branch diffusion model (DBDM)
The Communication Module (CM) in DBDM is based on the cross attention mechanism. Formally,
$\tilde{\bm{f}}^{h}=\operatorname{MLP}(\operatorname{Attn}({\bm{f}}^{h}{\mathbf{W}}_{Q},{\bm{f}}^{o}{\mathbf{W}}_{K},{\bm{f}}^{o}{\mathbf{W}}_{V})),$ (8)
$\tilde{\bm{f}}^{o}=\operatorname{MLP}(\operatorname{Attn}({\bm{f}}^{o}{\mathbf{W}}_{Q},{\bm{f}}^{h}{\mathbf{W}}_{K},{\bm{f}}^{h}{\mathbf{W}}_{V})),$ (9)
where $\operatorname{MLP}(\cdot)$ denotes fully-connected layers, $\operatorname{Attn}(\cdot)$ is the attention block [vaswani2017attention], and ${\mathbf{W}}_{Q},{\mathbf{W}}_{K},{\mathbf{W}}_{V}$ are learned projection matrices for query, key, and value, respectively.
The training objective of this full model is based on reconstruction loss
$\begin{split}\mathcal{L}_{hoi}=\mathbb{E}_{t\sim[1,T]}\|M_{\theta}(\bm{x}_{t},t,\bm{c})-\bm{x}_{0}\|^{2}_{2},\end{split}$ (10)
where $\bm{x}_{0}$ is the ground truth of the HOI sequence.
A.2 Affordance prediction diffusion model (APDM).
Model architecture. The affordance prediction diffusion model comprises eight Transformer layers for the encoder with a PointNet++ [qi2017pointnet++] to encode the object’s point clouds. The training objective of this diffusion model is also based on reconstruction loss
$\begin{split}\mathcal{L}_{aff}=\mathbb{E}_{t\sim[1,T]}\|{A}_{\theta}(\bm{y}_{t},t,\bm{p},\bm{d})-\bm{y}_{0}\|^{2}_{2},\end{split}$ (11)
where $\bm{y}_{0}$ is the ground-truth affordance data. $\bm{p}$ and $\bm{d}$ denote object point cloud and text description (prompt), respectively. $A_{\theta}$ represents the affordance prediction diffusion model.
Inferring object state with GPT-3.5-turbo in APDM. To infer the state of an object, we directly leverage the strong prior knowledge of large language models to derive the result. Specifically, we utilize the GPT-3.5-turbo [chatgpt] API by inputting specific instructions, allowing it to infer the result directly based on the input HOI text description. The prompt template for instruction is shown in Figure 6.
A.3 Affordance-guided interaction correction.
During the inference stage, it’s found that the predicted object contact positions may occasionally be inaccurately positioned, residing either inside or outside the object. To rectify this, we implement post-processing steps that replace these predicted contact points, denoted as $\bm{y}_{0}^{o}$, with their nearest neighbors from the object’s point clouds. This adjustment aims to enhance the accuracy of the updated contact points, aligning them more closely with their actual positions on the object’s surface. However, employing these updated contact points directly for contact constraints, particularly in the absence of detailed human shape information, introduces a new challenge. It can potentially lead to penetration issues within the contact area while reconstructing the human mesh in the final stage. To mitigate contact penetration, we adopt a method that recalculates points at a specified distance outward, perpendicular to the normal, originating from the object’s contact points. This process can formulated as: $\tilde{\bm{y}}{0}^{o}=\hat{\bm{y}}_{0}^{o}+v_{n}^{i}*d$, where $i\in\{1,2\}$ indicates the $i^{th}$ object contact points, $v_{n}^{i}$ denotes the normal vector at that point and $d=0.05$ is a contact distance threshold.
As for smoothness term, we formulate it as
$\begin{split}\small{G_{smo}=\sum_{l=1}^{L-1}\left\lVert{\bm{x}}_{0}^{o}(l+1)-{\bm{x}}_{0}^{o}(l)\right\rVert^{2},}\end{split}$ (12)
where ${\bm{x}}_{0}^{o}(l)$ is the predicted 6DoF pose of the object in the $l$-th frame.
Figure 6: Prompt template for inferring object state.
Algorithm 1 Affordance-guided Interaction Correction
1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$.
2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$
3:$K$ = 1
4:for all $t$ from $T$ to $1$ do
5:       ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$
6:       if $t=1$ then
7:             $K=100$
8:       end if
9:       for all $k$ from $K$ to $1$ do # Separately perturb
10:             ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$
11:       end for
12:       $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$
13:end for
14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Appendix B Implementation Details
Both our DBDM and APDM are built on the Transformer [vaswani2017attention] architecture. Similar to MDM [tevet2023human], we employ the CLIP model to encode text prompts, adhering to a classifier-free generation process. Our models are trained using PyTorch [paszke2019pytorch] on 1 NVIDIA A5000 GPU. We set control strength of guidance as $\tau_{1}$ = 1, $\tau_{2}$ = 100, and ${\Sigma}_{t}=min(\Sigma_{t},0.01)$. Both the DBDM and APDM are trained on the same data for 20k steps.
Both the DBDM and APDM architectures of HOI-Diff are based on Transformers with 4 attention heads, a latent dimension of 512, a dropout of 0.1, a feed-forward size of 1024, and the GeLU activation [hendrycks2016gaussian]. The number of learned parameters for each model is stated in Table 9.
Our training setting involves 20k iterations for the DBDM and 10k iterations for the APDM model. These iterations utilize a batch size of 32 and employ the AdamW optimizer [loshchilov2017decoupled] with a learning rate set at 10-4. We use $T$=1000 and $N$=500 diffusion steps in DBDM and APDM, respectively.
Appendix C Additional Details of Baselines
  • MDMfinetuned: We finetune MDM [tevet2023human] on BEHAVE dataset without considering the object motion.
  • MDM*: We extend the original feature dimensions of the input and output processing in MDM [tevet2023human] from $D^{h}$ to $D^{h}+D^{o}$, enabling support for HOIs sequences. The model is trained from scratch on BEHAVE dataset [bhatnagar22behave].
  • PriorMDM*: The proposed approach for dual-person motion generation employs paired fixed MDMs [tevet2023human] per individual to ensure uniformity within generated human motion distributions. This design leverages a singular ComMDM to coordinate between the two branches of fixed MDM instances, streamlining training and maintaining consistency across generated motions. Given that both branches are based on MDM that pretrained on human motion datasets, direct utilization of them for human-object interactions in our task is infeasible. We maintain one branch dedicated to humans, leveraging pre-trained weights, while adapting the input and output processing of another branch specifically for generating object motion. Following this, we fine-tune the human MDM branch while initiating the learning of object motion from scratch within the object branch. Eventually, we integrate ComMDM to facilitate communication and coordination between these distinct branches handling human and object interactions.
  • InterDiff: InterDiff [xu2023interdiff] is originally designed for a prediction task rather than text-driven HOIs generation. To tailor it to our task, we replace its Transformer encoder with a CLIP encoder and modify its feature dimensions of the input and output layers.
To ensure fair comparisons, all the above baselines as well as our own models are all trained on BEHAVE and OMOMO datasets for 20k steps.
Figure 7: Effect of different total numbers of perturbations in the whole denoising process. (a) Perturb one time in each denoising step (in total $T=1000$). (b) Perturb one time in first $T-1$ denoising steps, and repeatedly perturb 10 times in the final denoising step. (c) Perturb one time in first $T-1$ denoising steps, and repeatedly perturb 100 times in the final denoising step.
Appendix D Additional Details of Evaluation Metrics
For detailed information regarding metrics employed in human motion generation, including FID, R-Precision, and Diversity, we refer readers to [tevet2023human, Guo_2022_CVPR] for comprehensive understanding.
Contact Distance. Expanding on the concept of Contact Distance, we utilize the chamfer distance metric to quantify the closeness between human body joints and the object surface. This computation leverages ground-truth affordance data that includes human contact labels and object contact points,
$ContactDistance=\frac{1}{L}\sum_{l}^{L}{CD(\hat{\bm{x}}_{l}^{h},\hat{\bm{p}}_{l})},$ (13)
where $\hat{\bm{x}}_{l}^{h}$ represents two human contact joints at the $l$-th frame, indexed according to ground-truth contact labels. Additionally, $\hat{\bm{p}}_{l}$ denotes two object contact points derived from the object motion $\bm{x}^{o}_{l}$ at frame $l$, also indexed based on ground-truth information. $CD$ denotes the chamfer distance.
Penetration Score. We followed the  li2023controllable to compute the penetration score (Pene), each vertex of the body ($V_{i}$) is queried against the precomputed Signed Distance Field (SDF) of the object. This process yields a corresponding distance value for each vertex. The penetration score is then formalized as:
$Pene=\frac{1}{n}\sum_{i=1}^{n}{|min(d_{i},0)|},$ (14)
measured in centimeters (cm).